OpenStudy (anonymous):

f/g = ?/g^2 get a equivalent fraction

OpenStudy (anonymous):

are you satisfied with this ?: \( {f \over g } = { f.g \over g^2} \)

OpenStudy (anonymous):

that doesnt seem rightt.. idk

OpenStudy (anonymous):

f/g = ?/g^2 x = g^2 * f/ g = g* f